OpenStudy (96mertcank96):

what is one- half of 2^6 ? what is half of 2^6 ?

OpenStudy (anonymous):

1/6?

OpenStudy (anonymous):

one half of 2^6, you can write it as: \[\frac{2^6}{2}\] Right??

OpenStudy (anonymous):

2^6 is basically multiplying 2, 6 times.. Like: \[2^6 = 2 \times 2 \times 2 \times 2 \times 2 \times 2\]

OpenStudy (anonymous):

you can figure that out by reducing the powers to something easier numbers. \[2^6= 2^3*2^3\] can you work with that?

OpenStudy (96mertcank96):

yeah i could do that the difficult part is how the questions are worded i don't understand it.

OpenStudy (anonymous):

One half of 2^6 means : \[One \; \; Half = \frac{1}{2}\] So One half of 2^6 becomes : \[\frac{1}{2} \times 2^6 = ??\]

OpenStudy (96mertcank96):

1/2 x 2^6 = 32 ??????

OpenStudy (anonymous):

Not at all bad..

OpenStudy (tkhunny):

\(\dfrac{1}{2}\cdot 2^{6} = \dfrac{2^{6}}{2} = 2^{5} = 32\) There it is!

OpenStudy (anonymous):

And remember: \[\frac{x^a}{x^b} = x^{a - b}\]

OpenStudy (96mertcank96):

then what is half of 2^6?

OpenStudy (anonymous):

One and the same thing.. 7 days or a Week..

OpenStudy (96mertcank96):

So Same Answer?

OpenStudy (anonymous):

Figure it out yourself... :)

OpenStudy (tkhunny):

Rephrase: \(One\;Half\;of\;2^{6} = \dfrac{1}{2}\cdot 2^{6} = \dfrac{2^{6}}{2} = 2^{5} = 32\)

OpenStudy (anonymous):

\[One \; \; Half = \frac{1}{2} \quad \quad Half = \frac{1}{2}\]

OpenStudy (anonymous):

Two half will be: two multiplied by half.. Three half will be three multiplied by half.. \[Two \; \; Half = 2 \times \frac{1}{2} = 1\]

OpenStudy (tkhunny):

\(Two\;Quarters = \dfrac{1}{2}\\A\;Half = \dfrac{1}{2}\)

OpenStudy (anonymous):

\[A \; \; Quater = \frac{1}{4}\]

OpenStudy (96mertcank96):

yeah imma just put 2^5

OpenStudy (tkhunny):

\(One\;Halfling = Frodo\) - Just an example.

OpenStudy (anonymous):

Yeas you can...

OpenStudy (anonymous):

*Yes..